The monkey’s mother is half as old as the monkey will be when it is three times as old as its mother was when she was half as old as the monkey will be when it is twice as old as it is now.
The combined ages of the monkey and its mother are thirty years.

Answer by ankor@dixie-net.com(22740)

Let x = mother's age now. Let y = offspring's age now
:
The monkey’s mother is half as old as the monkey will be when it is three times as old as its mother was when she was half as old as the monkey will be when it is twice as old as it is now.
:
The last phrase:"when she was half as old as the monkey will be when it is twice as old as it is now." just means: .5(2y) = y
:
Replacing that phrase with y:
mother is half as old as the monkey will be when it is three times as old as its mother was y
:
x = .5(3y)
x = 1.5y
:
"The combined ages of the monkey and its mother are thirty years."
x + y = 30
1.5y + y = 30; substituted 1.5y for x
2.5y = 30
y = 30/2.5
y = 12 yrs old
Then
x = 18 yrs old
:
Check solution in:
mother is half as old as the monkey will be when it is three times as old as its mother was when y
18 = .5(3(12)
18 = .5(36)
:
I'm never sure of these tormentors, but no one seemed to want it, hence my effort. Hope it's useful to you.